OpenStudy (anonymous):

Solve the exponential equation. 9^8x = 27

OpenStudy (anonymous):

\[9^{8x}=27\]

OpenStudy (raden):

9^8x = 27 (3^2)^8x = 3^3 3^16x = 3^3 the bases already is same, now take the exponents then solve for x. 16x = 3 x = 3/16

OpenStudy (anonymous):

THANKSSS!(: thats what i thought! that is the final answer right? @RadEn

OpenStudy (raden):

yes, that is the final answer. you're welcome

OpenStudy (anonymous):

happy new year!(: @RadEn

OpenStudy (raden):

thanks. happy new year too :)